# Some remarks on Gorenstein projective precovers

### v2

@@ 4. Generalized Gorenstein flat modules
We will apply [15, Theorem 1.2]. Note that the hereditary and complete cotorsion pairs $([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m},\;[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ and $(\mathcal{P}(R)^{\wedge}_{m},[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}})$ are compatible, since from the proof of Lemma 4.13 we know that $\mathcal{P}(R)^{\wedge}_{m}=\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\cap[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. While the last assertion follows from [22, Theorem 6.21]. ∎
Finally, to obtain a further application of the theory developed so far, we will address a duality pair that we believe to be very interesting.
@@
Wakamatsu introduced in [24] and studied the named generalized tilting modules, usually called Wakamatsu tilting modules. Note that a bimodule ${}_{S}C_{R}$ is semidualizing if and only if it is Wakamatsu tilting [25, Corollary 3.2].
@@
Recently been shown in [19] that there are two duality pairs [^5] Note that in the reference [19] the bimodule is ${}_{R}C_{S}$, and in our text is ${}_{S}C_{R}$ as appear in [3] associated to the bimodule ${}_{S}C_{R}$. Rewritten for the semidualizing bimodule ${}_{S}C_{R}$ (this is; in [19] we change $R$ by $S$) as follows.
By [19, Theorem 3.3] in $\mathrm{Mod}(S^{\mathrm{op}})$, the pair
@@
is a symmetric duality pair, and by [19, Theorem 3.3] in $\mathrm{Mod}(R)$ the pair
@@
is a symmetric duality pair. Furthermore, the pairs $(\mathcal{A}_{C}(S^{\mathrm{op}}),\mathcal{B}_{C}(S))$ and $(\mathcal{A}_{C}(R),\mathcal{B}_{C}(R^{\mathrm{op}}))$ are perfect duality pairs, respectively by [19, Corollaries 3.4 & 3.6]. From the above we have the following result, whose proof follows from Proposition 4.9.
@@
Let ${}_{S}C_{R}$ be a semidualizing $(S,R)$-bimodule, and consider the associated Auslader classes $\mathcal{A}_{C}(R)$ and $\mathcal{A}_{C}(S^{\mathrm{op}})$. The following statements are true.
  • (a)
    If $R$ is left n-perfect then the class of Gorenstein $\mathcal{A}_{C}(R)$-projectives, denoted by $\mathcal{GP}_{\mathcal{A}_{C}}(R)$ is special precovering over $\mathrm{Mod}(R)$.
  • (b)
    If $S^{\mathrm{op}}$ is left n-perfect then the class of Goresntein $\mathcal{A}_{C}(S^{\mathrm{op}})$-projectives $\mathcal{GP}_{\mathcal{A}_{C}}(S^{\mathrm{op}})$ is special precovering over $\mathrm{Mod}(S^{\mathrm{op}})$.
  • (c)
    If $R$ is left n-perfect and $\mathcal{A}_{C}(R)\subseteq\mathcal{P}(R)^{\wedge}$, then the class $\mathcal{GP}(R)$ is special precovering on $\mathrm{Mod}(R)$
  • (d)
    If $S^{\mathrm{op}}$ is left n-perfect and $\mathcal{A}_{C}(S^{\mathrm{op}})\subseteq\mathcal{P}(S^{\mathrm{op}})^{\wedge}$, then the class $\mathcal{GP}(S^{\mathrm{op}})$ is special precovering on $\mathrm{Mod}(S^{\mathrm{op}})$.
@@
Funding The author was fully supported by a CONAHCyT posdoctoral fellowship at the Centro de Ciencias Matemáticas, UNAM.

### v3

@@ 3.1. Precovers from generalized Gorenstein $R$-modules
We will apply [21, Theorem 1.2]. Note that the hereditary and complete cotorsion pairs $([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m},\;[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ and $(\mathcal{P}(R)^{\wedge}_{m},[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}})$ are compatible, since from the proof of Lemma 3.15 we know that $\mathcal{P}(R)^{\wedge}_{m}=\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\cap[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. While the last assertion follows from [30, Theorem 6.21]. ∎
As stated in the introduction, we are interested in the contravariant finiteness of the class $\mathcal{GP}(R)^{<\infty}_{fin}$,  [^4] Note that $\mathcal{GP}(R)_{\infty}$ in the notation of [29] refers to $\mathcal{GP}(R)^{\wedge}=\mathcal{GP}(R)^{<\infty}$ in this paper. since this implies the contravariant finiteness of the class $\mathcal{P}(R)^{<\infty}_{fin}$ (such property implies the second finitistic dimension conjeture over an Artin algebra) over rings where $\mathcal{GP}(R)_{fin}$ is contravariantly finite. We will prove a kind of converse over $\mathrm{Mod}(R)$, in the generality of assuming that $\mathcal{GP}_{\mathcal{L}}$ is special precovering. We specify these ideas in the following result. But first we need to state a few notions adapted to our setting.
Definition 3.18.
[5, Definition 4.16] The finitistic $(\mathcal{P}(R),\mathcal{L})$-Gorenstein proyective dimension of $\mathrm{Mod}(R)$ is defined and denoted by
$\mathrm{FGPD}_{(\mathcal{P},\mathcal{L})}(R):=\sup\{\mathrm{Gpd}_{\mathcal{L}}(M):M\in\mathcal{GP}_{\mathcal{L}}^{\wedge}\}.$
where $\mathrm{Gpd}_{\mathcal{L}}(M):=\mathrm{resdim}_{\mathcal{GP}_{\mathcal{L}}}(M)$.
The finitistic projective dimension of $\mathrm{Mod}(R)$ is $\mathrm{FPD}(R):=\sup\{\mathrm{pd}(M):M\in\mathcal{P}(R)^{\wedge}\}$.
Note that the following result have a version over each case of Proposition 3.14.
Proposition 3.19.
Let $R$ be a ring and $\mathcal{L}\subseteq\mathrm{Mod}(R)$ such that $(\mathcal{P}(R),\mathcal{L})$ is a GP-admissible pair. If $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is a complete cotorsion pair and $\mathrm{FPD}(R)=t<\infty$ then $\mathcal{GP}_{\mathcal{L}}^{<\infty}$ is the left hand side of a complete and hereditary cotorsion pair.
Proof.
Indeed, since $\mathrm{FPD}(R)=t$, from [5, Theorem 4.23 (a)] we get that $\mathrm{FGPD}_{(\mathcal{P},\mathcal{L})}(R)=t<\infty$. This implies from Theorem 3.16 that $\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{t}=\mathcal{GP}_{\mathcal{L}}^{<\infty}$ is the left hand side of a complete and hereditary cotorsion pair. ∎
In the following, we compare the results obtained in this work with the second main result of P. Moradifar and J. Šaroch [29, (2.5) Theorem].
Remark 3.20.
Let us consider $\Lambda$ an Artin algebra. We know from [6, X, Theorem 2.4 (iv)] that $\mathcal{GP}(\Lambda)$ is the left hand side of a complete and hereditary cotorsion pair in $\mathrm{Mod}(\Lambda)$, so $\mathcal{GP}(\Lambda)$ is contravariantly finite in $\mathrm{Mod}(\Lambda)$. The following statements are true.
  • (i)
    For all $n\geq 0$, the class $\mathcal{GP}(\Lambda)^{\wedge}_{n}$ es contravariantly finite (compare with [29, (2.5) Theorem (i)]).
  • (ii)
    If the class $\mathcal{P}(\Lambda)^{<\infty}_{fin}$ of $\Lambda$-modules finitely generated of finite projective dimensiones contravariantly finite, by Huisgen-Zimmermann and Smalø [27], we get that $\mathrm{FPD}(\Lambda)<\infty$. Thus, from Proposition 3.19 we conclude that $\mathcal{GP}(\Lambda)^{\wedge}$ is contravariantly finite (compare with [29, (2.5) Theorem (ii)]).
Finally, to obtain a further application of the theory developed so far, we will address a duality pair from which we will be able to define a interesting class of Gorenstein $R$-modules on which we will apply our theory.
@@
Wakamatsu introduced in [32] and studied the named generalized tilting modules, usually called Wakamatsu tilting modules. Note that a bimodule $_{S}C_{R}$ is semidualizing if and only if it is Wakamatsu tilting [33, Corollary 3.2].
@@
Recently been shown in [26] that there are two duality pairs [^5] Note that in [26] the bimodule is $_{R}C_{S}$, and in our text is $_{S}C_{R}$ as appear in [25] associated to the bimodule $_{S}C_{R}$. Rewritten for the semidualizing bimodule $_{S}C_{R}$ (this is; in [26] we change $R$ by $S$) as follows.
By [26, Theorem 3.3] in $\mathrm{Mod}(S^{\mathrm{op}})$, the pair
@@
is a symmetric duality pair, and by [26, Theorem 3.3] in $\mathrm{Mod}(R)$ the pair
@@
is a symmetric duality pair. Furthermore, the pairs $(\mathcal{A}_{C}(S^{\mathrm{op}}),\mathcal{B}_{C}(S))$ and $(\mathcal{A}_{C}(R),\mathcal{B}_{C}(R^{\mathrm{op}}))$ are perfect duality pairs, respectively by [26, Corollaries 3.4 & 3.6]. From the above, we have the following result, whose proof follows from Proposition 3.12.
@@
Let $_{S}C_{R}$ be a semidualizing $(S,R)$-bimodule, and consider the associated Auslander class $\mathcal{A}_{C}(R)$. If $R$ is left n-perfect then the class $\mathcal{GP}_{\mathcal{A}_{C}}(R)$ of Gorenstein $\mathcal{A}_{C}(R)$-projectives is special precovering. If in addition $\mathcal{A}_{C}(R)\subseteq\mathcal{P}(R)^{\wedge}$, then the class $\mathcal{GP}(R)$ is also special precovering.
Finally in (ii) of the following result, we will make use of a different technique that can also be used on other versions of Gorenstein projective $R$-modules, we refer to a kind of “reduction”, as in [17]. We will see that as an advantage of this technique, we can dispense the condition of being n-perfect in the above Proposition. For do this we recall the notion of $C$-injective $R$-modules [25, Definition 5.1] , which is the class $\mathcal{I}_{C}(R)$ consisting by $R$-modules of the form $\mathrm{Hom}_{S}(C,I)$ where $I\in\mathcal{I}(S)$.
Theorem 3.22.
Let $_{S}C_{R}$ be a semidualizing $(S,R)$-bimodule. The class $\mathcal{GP}_{\mathcal{A}_{C}}(R)$ is special precovering under the following situations.
  • (i)
    When $\mathcal{A}_{C}(R)\subseteq\mathcal{GP}_{\mathcal{A}_{C}}(R)^{\wedge}$.
  • (ii)
    If $\mathrm{pd}(\mathcal{I}_{C}(R))\leq m$ and $\mathrm{id}(\mathcal{A}_{C}(R))<\infty$.
Proof.
(a) From [16, Proposition 3.9], we actually know that $\mathcal{GP}(R)\subseteq\mathcal{A}_{C}(R)$, in consequence $\mathcal{GP}_{\mathcal{A}_{C}}(R)\subseteq\mathcal{A}_{C}(R)$ since $\mathcal{P}(R)\subseteq\mathcal{A}_{C}(R)$ from [25, Theorem 6.2]. This give us the containment $\mathcal{GP}_{\mathcal{A}_{C}}(R)\subseteq\mathcal{A}_{C}(R)\subseteq\mathcal{GP}_{\mathcal{A}_{C}}(R)^{\wedge}$, which implies that all $N\in\mathcal{A}_{C}(R)$ possesses a special $\mathcal{GP}_{\mathcal{A}_{C}}(R)$-precover, for the above mentioned, Facts 3.9 and [5, Theorem 4.1 (a)]. Also was recently proved that for a general ring the pair $(\mathcal{A}_{C}(R),\mathcal{A}_{C}(R)^{\bot})$ is a perfect cotorsion pair, thus, by [17, Theorem 1] we get that the class $\mathcal{GP}_{\mathcal{A}_{C}}(R)$ is special precovering.
(b) Now let us suppose that $\mathrm{pd}(\mathcal{I}_{C}(R))\leq m$ and $\mathrm{id}(\mathcal{A}_{C}(R))<\infty$, we will prove under this conditions that $\mathcal{A}_{C}(R)\subseteq\mathcal{GP}_{\mathcal{A}_{C}}(R)^{\wedge}_{m}$, and so the result will follow from (a). To do this, take $M\in\mathcal{A}_{C}(R)$. By [16, Proposition 3.12] there exists an exact sequence
$0\to M\to U^{0}\to U^{1}\to\cdots$
where each $U^{i}\in\mathcal{I}_{C}(R)$.
From the inequality $\mathrm{pd}(\mathcal{I}_{C}(R))\leq m$ and [9, Ch. XVII, §1 Proposition 1.3] we can construct the following commutative diagram
[…]
such that $P_{i}^{j}\in\mathcal{P}(R)$ for all $i\in\{0,1,\dots,m\}$ and all $j\geq 0$. With $Q_{i}:=\mathrm{Ker}(P^{0}_{i}\to P^{1}_{i})\in\mathcal{P}(R)$ for all $i\in\{0,\dots,m-1\}$ and
$\Lambda:0\to Q\to P_{m}^{0}\to P_{m}^{1}\to\cdots$
$\Theta:0\to Q\to Q_{m-1}\to\cdots\to Q_{0}\to M\to 0$
exact complexes. Note that the complex $\Lambda$ can be completed on the left with a resolution by projective $R$-modules, and since $\mathrm{id}(\mathcal{A}_{C}(R))<\infty$, by [3, Lemma 3.6] such exact complex is $\mathrm{Hom}_{R}(-,\mathcal{A}_{C}(R))$-exact, from where $Q\in\mathcal{GP}_{\mathcal{A}_{C}}(R)$. But then, from the exact complex $\Theta$ we obtain that $\mathrm{resdim}_{\mathcal{GP}_{\mathcal{A}_{C}}(R)}(M)\leq m$, that is $M\in\mathcal{GP}_{\mathcal{A}_{C}}(R)^{\wedge}_{m}$. ∎
The previous result allows us to obtain another family of cotorsion pairs and a family Hovey triples similarly to the Corollary 3.17. By another hand, the following result address the conditions of when the class $\mathcal{GP}_{\mathcal{A}_{C}}(R)$ coincides with the classical $\mathcal{GP}(R)$.
Corollary 3.23.
Let $_{S}C_{R}$ be a semidualizing $(S,R)$-bimodule and assume that $\mathcal{A}_{C}(R)\subseteq\mathcal{P}(R)^{\wedge}$. Then, the class $\mathcal{GP}(R)$ is special precovering.
Proof.
Indeed, the condition $\mathcal{A}_{C}(R)\subseteq\mathcal{P}(R)^{\wedge}$ implies that $\mathcal{A}_{C}(R)\subseteq\mathcal{GP}_{\mathcal{A}_{C}}(R)^{\wedge}$, since by [25, Theorem 6.2] $\mathcal{P}(R)\subseteq\mathcal{A}_{C}(R)$. Thus, from Theorem 3.22 (a), the class $\mathcal{GP}_{\mathcal{A}_{C}}(R)$ is special precovering. Currently, we also have the containments $\mathcal{P}(R)\subseteq\mathcal{A}_{C}(R)\subseteq\mathcal{P}(R)^{\wedge}$, thus from [5, Proposition 6.7] we obtain the equality $\mathcal{GP}_{\mathcal{A}_{C}}(R)=\mathcal{GP}(R)$. ∎
@@
Funding The author was fully supported by a CONAHCyT (actually renamed SECIHTI) posdoctoral fellowship CVU 443002, managed by the Universidad Michoacana de San Nicolas de Hidalgo at the Centro de Ciencias Matemáticas, UNAM.
Declarations
Ethical Approval Not applicable.
Availability of data and materials Not applicable.
Competing interests The authors declare no competing interests.